OpenStudy (anonymous):

find the center and radius of the circle x^2 - 6x + y^2 - 4y + 4 = 0

OpenStudy (mertsj):

Complete the square on the x^2-6x part and factor the y^2-4y+4 part.

OpenStudy (mertsj):

What do you get?

OpenStudy (anonymous):

i know how to get the y^2 - 4y + 4 is (y-2)(y-2)

OpenStudy (mertsj):

so that would be (y-2)^2

OpenStudy (mertsj):

So now we have x^2-6x +(y-2)^2=0

OpenStudy (mertsj):

Do you know how to complete the square?

OpenStudy (anonymous):

i dont know how to complete the square of x^2 - 6x

OpenStudy (anonymous):

Find the center of this circle x^2-6x+y^2-4y+4=0 x^2-6x+9 + y^2-4y+4 = 9 (x-3)^2 + (y-2)^2 = 3^2 That's a circle of radius 3 with a center at (3,2) --------------- The other circle has the same center (concentric). Find its radius: r^2 = 3^2 + 2^2 r^2 = 13 -------- --> (x-3)^2 + (y-2)^2 = 13 or, x^2-6x+y^2-4y=0

OpenStudy (mertsj):